# Finding Visual Task Vectors

### v1

@@ 3.1 Scoring Activations
Low 3 (27, 5) 0.0281 -0.0280 15.4219
Note that since our end goal is to operate in a “zero-shot” setting (e.g, no input-output examples), we only consider the score and task vectors that correspond to the CLS token, the input query, and the output (present in the decoder only). Also, computing the token score and mean activations is very efficient as it simply requires applying the forward pass of a neural network across a batch without any access to a held-out validation set.
Importantly, while $\rho_{token}(\cdot)$ assigns high values to task vectors, it might also assign high values for other potential activations, for example, an activation that captures the color histogram can also vary across tasks (segmentation maps compared to natural images) but remain similar within each individual task. Therefore, we use this scoring mechanism mainly to build intuition about the existence of task vectors.
Finally, we compute the aggregated per-layer score:
$\rho_{layer}(l)=\sum_{i=(l,m,k)}\rho_{token}(i)$ (4)
Using this scoring function, we find that some activations capture “taskness” (see Figure 2) and cluster the data well with respect to tasks (see Table 1). We further elaborate on this analysis in Section 4 and Section 5. To find visual task vectors, we leverage a more general approach outlined in the following section.
3.2 Finding Visual Task Vectors via REINFORCE
Figure 2: Activation Scoring Analysis. Individual scores (${\rho_{token}}(i)$) aggregated per Attention Head (left). Individual token scores of specific heads are further visualized, along with the t-SNE [27] clustering of head activations of different tasks (right).
For every task $j$ we can apply the following procedure to find the task vectors. Recall that $\{\mu_{i,j}\}$ denote the mean activations and $F(\cdot)$ is a pretrained visual prompting model. Denote $\alpha_{i,j}\sim Bernoulli(\sigma(\theta_{ij}))$ as random variables that signify whether the mean task activation $\mu_{i,j}$ is a task vector and $\theta_{ij}$ is a learned weight followed by the sigmoid function to ensure it is in $[0,1]$.
Denote $z_{j}$ as the set of task vectors for task $j$: $z_{j}=\{\mu_{i,j}|\alpha_{i,j}=1\}$. Given pairs of input-output task demonstrations $x_{q},y_{q}\sim\mathbf{D_{task\_j}}$, we want to find the set of task vectors that minimizes the loss function:
$L(\theta)=E_{{z_{j}}\sim p_{\theta}}L(y_{q},F(x_{q}|z_{j}))$ (5)
This procedure is outlined for finding the task vectors placements for individual tasks, however it is possible to find placements that generalize for multiple tasks by optimizing over examples across datasets and we refer this as “multi-task” patching. Additionally, while above we search for all token activations, it is possible to apply the search in different levels of granularity. For example, by patching groups of tokens from the same quadrant, patching all the tokens in an attention head, or patching the entire layer. We discuss these design choices in the next section.
4 Experiments
Our experiments aim to explore whether activation patching of task vectors in zero-shot environments can cause the model to execute the desired visual task, hopefully performing similarly–or better–to the original model’s one-shot in-context-learning performance. In this section, we describe the implementation details of MAE-VQGAN, the prompting schemes used, the details of our implementation, the baselines used for comparison, the visual tasks of interest, and the experiments conducted.
4.1 Implementation Details
MAE-VQGAN [4]. An MAE  [16] with a ViT-L [9] backbone, which has $24$ encoder blocks and $8$ decoder blocks, with $16$ attention heads in every layer and a patch size of $1616x$ 16 and input resolution of $224224x$ 224. The decoder predicts a distribution over a VQGAN [10] codebook to output images with better visual quality. We used the pretrained checkpoint from [4] trained over the Computer Vision Figures [4] dataset and ImageNet [35].
Zero-shot Prompting. Similarly to one-shot with the key difference that only the query is embedded into the grid image, in the same bottom left quadrant position as in the one-shot setting. The model is then used to reconstruct only the part that corresponds to the output. Specifically, we patchify the query image at a resolution of 112x112 and apply to it the positional encodings of the bottom left quadrant. The patches are then fed into the encoder and are processed by the decoder together with the mask tokens that correspond to the bottom right quadrant to obtain the result. Note that in the zero-shot setting we also patch intermediate activations by task vectors.
Finding Task Vectors. We use REINFORCE  [43] to select where to patch. We initialize $\theta_{ij}$ with $-1$. Each patch position modeled as a random variable corresponds to either the CLS token, or a grouping of the image patch tokens based on their spatial positioning: the bottom left quadrant (the image patch tokens that correspond to the query), and bottom right quadrant (the MASK tokens). In other words, we restrict the granularity of the patching procedure; instead of having an individual random variable parameter for each token position $i$ we reduce the dimensionality of the search space by grouping the tokens into the three categories mentioned. Upon each iteration of REINFORCE, we sample $32$ times from the Bernoulli distribution for each image (fixed at $10$ for all experiments) and perform a patching procedure into the positions sampled with a value equal to $1$. This results in a total of $320$ executions of zero-shot MAE-VQGAN per iteration. Then, we optimize the Bernoulli parameters as outlined in 3.2 with Adam [20] using a learning rate of $0.1$. We execute the algorithm for a total of $600$ steps and select the checkpoint at intervals of 50 steps with the best evaluation on a held-out test set.
4.2 Activation Scoring Analysis
Here our goal is to evaluate in isolation the Activation Scoring step (outlined in Section 3.1), specifically whether high scoring activations indeed correspond to Task Vectors.
Collecting Activations. As outlined in Section 3.1, our first step towards computing activation scores is to run the forward pass of the model in a one-shot setting to collect activations across different tasks. Specifically, we use $100$ prompts and queries from Pascal 5i [36] training set, ensuring that the one-shot performance across all tasks works reasonably well. During the forward pass, we save the activations for every task $j$ and $i=(l,m,k)$, which corresponds to the intermediate activation of the $k^{th}$ token in the $l^{th}$ block after the $m^{th}$ attention head. Afterwards, we compute the mean activation ${\mu_{i,j}}$ and score $\rho_{token}(i)$.
4.3 Downstream Tasks
Dataset. We use the Pascal-5i [36] dataset, which is comprised of 4 different image splits where every split contains between $346$ and $725$ images and associated segmentation masks. For evaluation, we sample $1000$ prompt-query pairs from each split in the validation set. For Activation Scoring and the various methods to find Task Vectors we only use examples from the training set.
Foreground Segmentation. For model evaluations, we use the segmentation masks included in the Pascal-5i [36] dataset, and report the mean IOU (mIOU) metric across the four splits.
Low Light Enhancement. To obtain the input-output data, given a Pascal-5i [36] image, we multiply the color channels by a factor of $0.5$ and define the result as the input and the original image as the output. We report the Mean Squared Error (MSE) metric of the prediction with the label.
Inpainting. To obtain input-output pairs, we randomly mask a square region in height/width equal to $25\%$ of an image (resulting in a black square of 1/8 the area) and define this as the input, while using the original image as the output. For evaluation, we report the MSE metric.
4.4 Ablations
In this section we describe the set of experiments conducted to ascertain the particular implementation details of our REINFORCE [43] method, validating our design choices.
Task Vectors Location in Encoder vs. Decoder. We hypothesize that task implementation occurs in a distributed fashion across multiple parts of the network, necessitating interventions in both the encoder and decoder to induce task implementation in the zero-shot scenario. Hence, we test this by executing our method by restricting interventions to the encoder only, the decoder only, and allowing for interventions throughout the whole network. We report the mIoU on the four splits for the Segmentation task, seeking to find if interventions in both parts of the model are required for appropriate task implementation.
Patching Granularity. We explore different intervention granularities by mapping each token position to a group thus decreasing the dimensionality at which we perform interventions. With this in mind, we can further reduce the size of the search space of the optimization by grouping potential patching positions $i=(l,m,k)$ into spatial quadrants or individual attention heads. We seek to find the optimal granularity at which we maintain precise interventions but manage to reduce the search space nonetheless. We execute our method with three granularity levels (individual tokens, quadrants, attention heads) and report the performance on the four tasks.
4.5 Zero-shot Task Vector Patching
We seek to explore whether zero-shot task vector patching can provide task implementation performance similar to one-shot prompting. We perform the following experiments to investigate this:
Multi-task. We follow the same procedure but in a multi-task scenario where we select two images and perform our method alongside the CMA and GRS baselines, evaluating the loss across all tasks jointly. We normalize the loss per task in order to account for task-specific differences to ensure no single task dominates the search, and all tasks are weighted equally. Finally, we introduce the identity-copy task to keep the batch size at 10 for equivalent comparisons.
MAE-VQGAN. We compare to the original model’s one-shot performance.
GRS across Random K Layers. We execute the search algorithm constraining it to a random set of $k$ layers instead of ranking them by their score. This serves to validate the need of the proposed scoring method.
We report the results across the four splits for the four tasks, and compare all variants to the original MAE-VQGAN model and the random baselines.
5 Results
5.1 Activation Scoring Analysis
We compute ${\rho_{token}}(i)$ and display it aggregated per head on a heatmap where the y-axis is the layer and the x-axis is the attention head index (see Figure 2). This heatmap showcases certain heads that stand out, suggesting that these heads may hold task vectors. From this heatmap we then select the top two heads ranked by score which are at position $(26,3)$ and $(11,3)$; we also select a lower-ranked head at position $(27,5)$. For each of these three heads, we visualize the clustering of its activations, and the individual Activation Score per token. Both are placed to the right side of the heatmap.
Clustering Visualization. We observe a clear clustering based on task in heads ranked highly–head (26,3) for instance–by our scoring methodology, and what appears to be many small clusters for the low-ranked head which we hypothesize are clustering by the semantics of the input prompt-query pair–head (27,5). To perform the dimensionality reduction we vertically stack the activations of a particular head across token positions and project them onto 2D with t-SNE [27] coloring by task.
Score-per-token Heatmap. We display the un-aggregated ${\rho_{token}}(i)$ values for each token re-arranged to convey spatial positioning (equivalent to the 2x2 prompting grid). That is, for the heads of interest, we report the ${\rho_{token}}(i)$ values that the particular individual head encapsulates–displayed on a heatmap. We place the CLS token as the lone square on the top left corner; following this we have the values corresponding to the four different quadrants $(x_{s},y_{s},x_{q},y_{q})$. It is important to note that in the encoder there are no tokens that correspond to the $y_{q}$ (bottom right) quadrant as these are incorporated as mask tokens after the encoder; hence, for visualization purposes we display their value to be marked as X. We observe that within a single head different tokens can take a wide range of values. Interestingly, there appears to be a consistency in values across the tokens in particular quadrants which motivated the decision to group token patching by quadrants.
Quantitative Clustering Analysis. To further support the observation of high and low-quality clustering based off of the scoring value, we report the Silhouette Score and Davies-Bouldin score for the four highest-ranked heads, 5 randomly sampled heads, and the three lowest-ranked heads (see Table 1). We observe that attention heads scored highly by our methodology display high quality clustering scores from the Silhouette Score and Devies-Bouldin Score and vice-versa for those scored lowly. Finally, the randomly selected attention heads received intermediate scores.
5.2 Zero-shot Task Vector Patching
We now share the performance of our method, both task-specific and multi-task, in comparison to the original model and the following baselines: Causal Mediation Analysis, Greedy Random Search, Random Quadrants, Random K Layers, and Top Quadrants. Surprisingly, by performing these interventions with task vectors we are able to implement visual ICL tasks with better performance than the original model. We observe in the following Table  2 (which displays the mean and variance across the 4 splits, or Table 5 in the Supplementary Materials which contains the individual split evaluations), that our task-specific models beat the original MAE-VQGAN across all tasks. Furthermore, the GRS task-specific models beat the original MAE-VQGAN in some tasks. We also observe that our multi-task models beat the original MAE-VQGAN in all tasks except Segmentation where performance is met, and GRS multi-task performs comparably to the original. CMA performs better than the random baselines but does not reach the high performance of our method, or the GRS method nonetheless.
Furthermore, we provide validation that ranking layers by their $\rho_{layer}(l)$ and selecting the top-k is instrumental to supporting the performance of the greedy random search by comparing to randomly selecting k layers. We see that random K layers does not perform well, except for colorization where it is similar to GRS. Moreover, if we simply patch into the top quadrants without the GRS we get poor performance.
@@ 5.2 Zero-shot Task Vector Patching
We now share the qualitative results for Segmentation, Lowlight Enhancement, and In-painting. First, we visualize the task-specific models of our methodology alongside the original MAE-VQGAN, and the CMA and GRS baselines (see Figure 3). Secondly, we visualize the task-specific and multi-task variants of our methodology in comparison to CMA and GRS(see Figure 4)
@@
5.3 Ablations
Task Vectors Location in Encoder vs. Decoder. We report our results on isolating the set of possible interventions to the encoder only, decoder only, in contrast to allowing interventions throughout the whole network. We can observe that in-context task learning builds upon both model components. The decoder, however, is more important. It is clear that intervening in both components is crucial for task implementation as we hypothesize that it is computed in a distributed fashion with cascading higher-order effects through the network where early interventions have strong downstream effects (see Table 3).
Segmentation $\uparrow$
Model Split 0 Split 1 Split 2 Split 3
Encoder (Task-specific) 0.09 0.14 0.14 0.13
Decoder (Task-specific) 0.32 0.34 0.29 0.29
Both (Task-specific) 0.35 0.35 0.31 0.29
Patching Granularity. Motivated by the emergence of quadrants in the per-token scoring visualization, we explore the optimal granularity at which to group the tokens to reduce the dimensionality of the search space. In the tasks of Segmentation and Colorization, grouping by quadrants results in better performance; whereas in the tasks of Lowlight Enhancement and In-painting, maintaining full patching granularity at the token level results in better performance (see Table 4 which displays the mean and variance across the 4 splits, or Table 6 in the Supplementary Materials which contains the individual split evaluations).
@@ 6 Limitations
Our focus was in identifying Task Vectors, but we hypothesize that there might exist other important vector-types such as image structure activations that capture locations, input-output placements, and left-to-right ordering. During the optimization with REINFORCE [43] we evaluate the performance using the MSE metric for all tasks (except Segmentation which uses mIoU) after decoding the predicted VQGAN tokens. Instead, it may be possible to directly evaluate the model in VQGAN token space using cross entropy loss which may lead to more accurate results. We leave these extensions for future works.
@@ 7 Conclusion
In this work we explore the internal mechanisms of visual in-context learning and devise an algorithm to identify Task Vectors, activations present in transformers that can replace the in-context examples to guide the model into performing a specific task. We confirm our approach by adapting MAE-VQGAN to perform specific tasks in a zero-shot fashion by patching the Task Vector identified. We find that different than in NLP, in computer vision Task Vectors are distributed throughout the network’s encoder and decoder.
Acknowledgements: This project has received funding from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (grant ERC HOLI 819080). Prof. Darrell’s group was supported in part by DoD including DARPA’s LwLL and/or SemaFor programs, as well as BAIR’s industrial alliance programs.
@@ 9 More Qualitative Examples
We provide a wider selection of examples comparing our zero-shot task vector patching methodology in comparison to 1) the original one-shot MAE-VQGAN, CMA, and GRS, and 2) our ablations. Alongside each figure, we accompany it with an according analysis.
@@ 10.1 Selecting Task Vectors via Greedy Random Search
Input. The mean activations $\{\mu_{i,j}\}$, scoring function $\rho_{layer}(\cdot)$, pretrained visual prompting model $F(\cdot)$, and an evaluation set.
@@
Greedy Search. Iterate over $l$ in the top $k$ scoring layers sorted by $\rho_{layer}(l)$ from high to low, pick activation $i=(l,m,k)$, flip the value $\alpha_{i,j}$ and evaluate the validation score for $F(\cdot|z_{j})$. After having evaluated each flip of the $\alpha_{i,j}$ on the particular layer $l$ we keep the $\alpha_{i,j}$ which performed best or continue to the next layer if the performance did not improve.
@@ 10.4 GRS Baseline Qualitative Comparisons
We provide a wider selection of examples comparing our GRS zero-shot task vector patching methodology in comparison to 1) a selection of baselines, and 2) our ablations. Alongside each figure, we accompany it with an according analysis.
@@
In the Figures 12, 13, and 14 we compare our GRS task-specific method with a handful of baselines defined in section 4.5. We have abbreviated Top Quadrants as Top Q, Random K Layers as Random Layers, and Random Quadrants as Random Q. It is clear that Top Quadrants struggles to output coherent completions. We believe this to be because of the need of patching into positions of different purposes other than task implementation such as positions that encode the input-output structure that a one-shot prompt provides. Further opportunities for exploration could include other scoring terms that take into account structural information provided by different prompt orientations. Random K Layers performs surprisingly well due to the efficiency of the Greedy Random Search but nonetheless does not reach the performance of using our scoring mechanism to select the top $K$ layers. Finally Random Quadrants struggles to complete coherent results.

### v2

@@ 5.1 Activation Scoring Analysis
5.2 Task Vector Patching
We present the performance of our task-specific and multi-task methods compared to the original model and baselines. Task Vector interventions improve visual ICL task performance over the original model. Table 2 shows mean and variance across 4 splits (full results in Table 5, Supplementary Materials). Our task-specific models outperform MAE-VQGAN in all tasks, with GRS models surpassing it in some. Our multi-task models excel in all tasks except Segmentation, where they match MAE-VQGAN. GRS multi-task matches the original model, while CMA improves upon random baselines but falls short of our method and GRS.
@@ 5.2 Task Vector Patching
We validate the importance of ranking layers by $\rho_{layer}(l)$ (Section 11.1) and selecting top-k for greedy random search performance. Random K layers generally underperform, except in colorization. Simply patching top quadrants without GRS also yields poor results. Qualitative results for Segmentation, Lowlight Enhancement, and In-painting are presented. We compare task-specific models of our method with original MAE-VQGAN, CMA, and GRS baselines (Figure 3), and visualize task-specific and multi-task variants against CMA and GRS (Figure 4).
5.3 Ablations
Task Vectors Location in Encoder vs. Decoder. We compare interventions isolated to the encoder, decoder, and throughout the whole network. Results show that in-context task learning utilizes both components, with the decoder playing a more crucial role. Intervening in both components is essential for task implementation, supporting our hypothesis of distributed computation with cascading effects throughout the network (see Table 4).
Patching Granularity. Inspired by quadrant patterns in per-token scoring, we explore optimal token grouping to reduce search space dimensionality. Quadrant grouping improves performance for Segmentation and Colorization, while token-level granularity is better for Lowlight Enhancement and In-painting (see Table 3 for mean and variance across 4 splits, and Table 6 in Supplementary Materials for individual split evaluations).
@@ 5.3 Ablations
Segmentation $\uparrow$
Model Split 0 Split 1 Split 2 Split 3
Encoder (Task-specific) 0.09 0.14 0.14 0.13
Decoder (Task-specific) 0.32 0.34 0.29 0.29
Both (Task-specific) 0.35 0.35 0.31 0.29
@@ 6 Limitations
While we focused on Task Vectors, other important vector-types might exist, for example, vectors capturing image structure and ordering. We evaluated performance using MSE (except mIoU for Segmentation) after decoding VQGAN tokens. Future work could explore direct evaluation in VQGAN token space using cross entropy loss for potentially more accurate results.
@@ 7 Conclusion
In this work we explore the internal mechanisms of visual in-context learning and devise an algorithm to identify Task Vectors, activations present in transformers that can replace the in-context examples to guide the model into performing a specific task. We confirm our approach by adapting MAE-VQGAN to perform tasks without including the ICL demonstration in the prompt by patching the Task Vectors identified. We find that different than in NLP, in computer vision Task Vectors are distributed throughout the network’s encoder and decoder.
Acknowledgements: This project has received funding from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (grant ERC HOLI 819080). Prof. Darrell’s group was supported in part by DoD including DARPA’s LwLL and/or SemaFor programs, as well as BAIR’s industrial alliance programs. This work was completed in partial fulfillment for the Ph.D degree of the last author.
@@ 8 Full Results
9 Additional Explorations
We include a handful of additional experiments that serve as potential future avenues for research.
Finding Task Vectors in Other Architectures. We evaluate our method to find Task Vectors on Llama2 7B, an autoregressive decoder-only architecture on NLP tasks and compare the Top-1 accuracy to previous reported results by Todd et al. [38]. Our method performs better than previous approaches, beating 10-shot in 2 tasks while reducing FLOPs by 92.5% (Table 7).
Table 7: Comparison of Task Vector methods on Llama2 7B for various NLP tasks
Method Landmark to Country Present to Past Country to Capital
0-shot 0.0 0.084 0.047
2-shot 0.868 0.95 0.92
10-shot 0.88 0.967 0.951
0-shot + CMA 0.691 0.88 0.825
0-shot+ Ours 0.8629 0.983 0.9524
Out of Domain Evaluation. We evaluate the performance of in-painting and low-light enhancement task-specific models on x-ray images. The qualitative results (Figure 5) show proper task implementation. The resulting x-ray images are slightly more blurry, which is due to a limitation in the underlying VQGAN tokenizer, previously reported in Bar et al. [4].
Figure 5: Out of domain evaluation on x-ray images.
Task Vector Arithmetic. We explore whether vector arithmetic can be performed on the task representations to create new tasks by composition. By utilizing the multitask patching positions, we can compose the tasks of in-painting and segmentation by combining their individual Task Vectors as follows: $combined=inpaint+segment-identity$ similarly to Todd et al. [38], where identity is the identity mapping task described in the paper. We compute these new mean activations and patch them into the previously determined multitask positions, and evaluate using as input a masked image and its corresponding segmentation. Figure 6 shows that the tasks are indeed composable, where $combined$ performs qualitatively better than $segment$ which suffers from holes.
Figure 6: Task vector arithmetic results.
One-shot Task Vectors. We investigate whether 1-shot performance can be improved with Task Vectors. Table 8 shows that additional in-context examples (or shots) are not necessary when using Task Vectors. Intuitively, Task Vectors already lead to improved performance and additional in-context learning examples do not add more information.
Table 8: Comparison of no ICL demonstration and one-shot performance with Task Vectors (TV)
Method Segm LowLight Inpaint Colorize
mIoU MSE MSE MSE
Original one-shot 0.338 0.6 0.55 0.618
No ICL + TV 0.353 0.458 0.480 0.453
One-shot + TV 0.346 0.496 0.454 0.490
@@ 10 More Qualitative Examples
We provide a wider selection of examples comparing our task vector patching methodology in comparison to 1) the original one-shot MAE-VQGAN, CMA, and GRS, and 2) our ablations. Alongside each figure, we accompany it with an according analysis.
@@ 11.1 Selecting Task Vectors via Greedy Random Search
Input. The mean activations $\{\mu_{i,j}\}$, pretrained visual prompting model $F(\cdot)$, an evaluation set, and the aggregated per-layer score defined as $\rho_{layer}(l)$:
$\rho_{layer}(l)$ $=\sum_{m,k}\rho_{token}(i=(l,m,k))$ (6)
@@
Greedy Search. Iterate over $l$ in the top $k$ scoring layers sorted by $\rho_{layer}(l)$ from high to low, pick activation $i=(l,m,k)$, flip the value $\alpha_{i,j}$ and evaluate the validation score for $F(\cdot|z_{j})$. After having evaluated each flip of the $\alpha_{i,j}$ on the particular layer $l$ we keep the $\alpha_{i,j}$ which performed best (based on a certain evaluation function) or continue to the next layer if the performance did not improve.
@@ 11.4 GRS Baseline Qualitative Comparisons
We provide a wider selection of examples comparing our GRS task vector patching methodology in comparison to 1) a selection of baselines, and 2) our ablations. Alongside each figure, we accompany it with an according analysis.
@@
In the Figures 14, 15, and 16 we compare our GRS task-specific method with a handful of baselines defined in section 4.6. We have abbreviated Top Quadrants as Top Q, Random K Layers as Random Layers, and Random Quadrants as Random Q. It is clear that Top Quadrants struggles to output coherent completions. We believe this to be because of the need of patching into positions of different purposes other than task implementation such as positions that encode the input-output structure that a one-shot prompt provides. Further opportunities for exploration could include other scoring terms that take into account structural information provided by different prompt orientations. Random K Layers performs surprisingly well due to the efficiency of the Greedy Random Search but nonetheless does not reach the performance of using our scoring mechanism to select the top $K$ layers. Finally Random Quadrants struggles to complete coherent results.